The curve has an equation y = e^x. Compute the area bounded by the


curve from x = 0 to x = 1.

1
Expert's answer
2022-06-08T06:18:28-0400

The area bounded by the curve "y=f(x)" from "x = a" to "x = b" can be found out by integration:

"A=\int_a^bf(x)dx."





We have:

"A=\int_0^1e^xdx=e^x |_0^1=e^1-e^0=e-1."


Answer: "e-1."